T1BINARY(1) T1BINARY(1) NAME t1binary−converttype-1fontfromASCIItobinary SYNOPSIS t1binary[−llength][input[output]] DESCRIPTION t1binaryconvertsAdobeType1fontprogramsinPFA(hexadecimal)formattoPFB(binary) format.Ifthefileoutputisnotspecifiedoutputgoestothestandardoutput.Ifthefileinputis notspecifiedinputcomesfromthestandardinput. OPTIONS −−block−length=length,−llength SetthemaximumlengthofsectionsinPFBoutputtolength.Thedefaultlengthisaslarge asmemoryallows. SEEALSO t1ascii(1),t1unmac(1),t1mac(1),t1disasm(1),t1asm(1) AdobeType1FontFormat AUTHORS LeeHetherington(ilh@lcs.mit.edu) EddieKohler(eddietwo@lcs.mit.edu) PortedtoMicrosoftC/C++CompilerandMS-DOSoperatingsystembyKai-UweHerbing(herb-ing@netmbx.netmbx.de). 5Mar1998 1